OpenStudy (anonymous):

Find the mean of the following group of numbers: 6.1, 5.9, 4.2, 5.9, and 10.7. Round your answer to the nearest tenth. 6.6 6.5 10.7 4.2

OpenStudy (anonymous):

mean =sum/no. of term =(6.1+5.9+4.2+5.9+10.7)/5 =6.56 =6.6 ( round to the nearest tenth)